Q7.The grades in a statistics course for a particular semester were as follows

Grade A B C D F
f 14 18 32 20 16

significance, that the distribution of grades is uniform.
1
Expert's answer
2019-07-15T10:19:46-0400

"H_0:" The distribution of grades is uniform.


"H_1:" The distribution of grades is not uniform.


Significance level: "\alpha=0.05"

n = number of observations = 14+18+32+20+16 = 100

k = number of intervals = 5


Test statistics:


"\chi^2=\sum{(f_0-f_e)^2\over f_e} \ with \ v=5-1=4, {n \over k}={100 \over 5}=20"


Critical Regions

The critical region is given by


"\chi^2>\chi_{\alpha, v}^2,\ \chi^2>\chi_{0.05, 4}^2, \ \chi^2>9.488"

Computation


"\chi^2=\sum{(f_0-f_e)^2\over f_e} ={(14-20)^2\over 20}+{(18-20)^2\over 20}+{(32-20)^2\over 20}+"

"+{(20-20)^2\over 20}+{(16-20)^2\over 20}=10"

Statistical decision

Since "10>9.488," then we reject "H_0."


Conclusion

The distribution of grades is not uniform.